OpenStudy (anonymous):

find the determinant of the matrix: (3x3) [x 0 -4] [3 x x^2] = ? [-3 x 1]

OpenStudy (anonymous):

\[D=\begin{vmatrix}x&0&-4\\ 3&x&x^2\\-3&x&1\end{vmatrix}\] Using a cofactor expansion along the first row, you have \[\begin{align*}D&=x\begin{vmatrix}x&x^2\\x&1\end{vmatrix}-0\begin{vmatrix}3&x^2\\-3&1\end{vmatrix}-4\begin{vmatrix}3&x\\-3&x\end{vmatrix}\\\\\\ &=x\begin{vmatrix}x&x^2\\x&1\end{vmatrix}-4\begin{vmatrix}3&x\\-3&x\end{vmatrix}\end{align*}\] Finding the determinant of a 2x2 is easy. Can you finish the work?

OpenStudy (anonymous):

Yeah, thanks

OpenStudy (anonymous):

yw